OpenStudy (anonymous):

Write two equations in standard form that are equivalent to the given equation 8x-14y=2

OpenStudy (retireed):

To me equivalent means equal but different. Kind of confusing huh? Margarine is used equivalently as butter, but different. If I divide or multiple every term by a number and I mean every term by the same number, is the equation equivalent but different? x | 2y = 10x + 8 | y = 5x + 4 --------|---------------|--------------- 0 | 2y = 8 y = 8/2 | y = 4 1 | y = 18/2 | y = 9 2 | y = 28/2 | y = 14 If you continue this you will see that for any x the two equivalent equation equal the same y. You should be able to write an infinite number of equations equivalent to 8x -14y = 2

OpenStudy (dayakar):

8x-14y=2----(1) multiply or divide the equation with any real number u can get equivalent equation

OpenStudy (dayakar):

example divide equation (1) by 2 u can get 4x-7y =1

OpenStudy (denisaboichuk):

11